# Influence of packing protocol on fractal exponents in dense polydisperse packings

### v1

@@ Abstract
We study fractal properties of a system of densely and randomly packed disks, obeying a power-law distribution of radii, which is generated by using various protocols: Delaunay triangulation (DT) with both zero and periodic boundary conditions and the constant pressure protocol with periodic boundary conditions. The power-law exponents of the mass-radius relation and structure factor are obtained numerically for various values of the size ratio of the distribution, defined as the largest-to-smallest radius ratio. It is shown that the size ratio is an important control parameter responsible for the consistency of the fractal properties of the system: the greater the ratio, the less the finite size effects are pronounced and the better the agreement between the exponents. For the DT protocol, the exponents of the mass-radius relation, structure factor, and power-law distribution coincide even at moderate values of the size ratio. By contrast, for the constant-pressure protocol, all three exponents are found to be different for both moderate (around 300) and large (around 1500) size ratios, which might indicate a biased rather than random spatial distribution of the disks. Nevertheless, there is a tendency for the exponents to converge as the size ratio increases, suggesting that all the exponents become equal in the limit of infinite size ratio.
dense random packings; jammed packings; constant pressure protocol; power-law polydispersity; fractals; structure factor; mass-radius relation.
@@ I Introduction
Compact packings with power-law size distributions are of great interest for various applications in science and engineering, in particular for understanding the structural properties of materials [1]. This understanding is critical for applications ranging from ultra-high performance concrete [2] to biological systems [3] and soils [4], as it enables the design and optimization of materials with desired mechanical properties and stability. The spatial correlations in such systems exhibit fractal-like behaviour [5, 6, 7, 8] and the description of the emerging fractal properties is one of the most interesting and intriguing problems.
In numerical simulations, dense packing can be achieved by various methods (see [1] for a review), including random successive addition (RSA) [9, 10]. In RSA, particles are added one at a time in random positions. If a particle overlaps with existing ones, it is discarded. This process continues until no more particles can be added without overlap. Recently, the authors suggested [8] Delaunay triangulation (DT) protocol as a modification of RSA. DT creates a network of triangles connecting the centers of packed particles, which optimizes the process of finding empty space for newly added particles.
Another method of dense packings is the constant pressure (CP) protocol, which uses LAMMPS (large-scale atomic/molecular massively parallel simulator) with GRANULAR package [11]. It allows the modeling of granular materials under controlled pressure conditions. Constant external pressure slowly compresses an initially dilute configuration of particles, leading to compact jammed packing.
It was shown with RSA [6] and DT [8] protocols that dense random packings of disks with a power-law size distribution exhibit fractal-like properties with the fractal dimension $D_{\mathrm{f}}$ being equal to the exponent of the power-law size distribution $D$. On the other hand, Monti et al. [7] obtained jammed packings of the same distribution with CP protocol and argued that the packings, although exhibiting fractal properties, nevertheless have fractal dimensions different from the exponent of the power-law size distribution. The authors of the paper [7] concluded that the fractal exponent depends on the packing protocol.
@@
As shown in the previous paper [6], the fractal range of densely packed disks or spheres with a power-law size distribution is determined by their largest-to-smallest size ratio, and this result is still valid in the thermodynamic limit [8]. In the paper [7], the ratio was taken in two dimensions from $100$ to $300$ for various exponents $D$ of the distribution [^1] Note that in the paper [7], the notation $\beta=D+1$ is used.. It is shown below in Sec. IV that these ratios are not large enough to obtain the correct value of the fractal dimension of the system. To this end, the data of jammed packings with CP protocol of Ref. [7] are reproduced for a specific value of the exponent $D=1.5$ and the size ratio $292$. We calculate both structure factor and mass-radius relation and obtain the discrepancy between the exponents $\alpha$ [$S(q)\sim 1/q^{\alpha}$] and $D_{\mathrm{f}}$ [$M(r)\sim r^{D_{\mathrm{f}}}$] in the corresponding fractal ranges: $\alpha\simeq 1.31$ and $D_{\mathrm{f}}\simeq 1.419$, see Fig. 2 and Table 1 below. This discrepancy is due to the insufficient size ratio.
Increasing the size ratio to $R/a=1575$ and keeping the other parameters unchanged (see Sec. IV.2 below), we obtain $\alpha\simeq 1.41$ and ${D_{\mathrm{f}}}\simeq 1.441$, which are much closer to each other, although its value is still different from $D=1.5$ predicted by our model of dense random packings [6, 8]. Nevertheless, this result suggests that the both exponents slowly converge to $D$ as the size ratio $R/a$ increases.
This paper is organized as follows. In Sec. II, we explain in detail a set of disk to be packed and packing methods. In the next section, the fractal properties of dense random packing with DT protocol is studied for various values of the size ratio. In Sec. IV, the results of packing with CP protocol are represented. In Conclusions, we summarize the main results and outline prospects for future research.
@@ II A set of disks to be packed and methods
Figure 1: Comparison of spatial distributions of $N=125000$ disks obtained from different protocols: (a) DT with zero boundary conditions (packing fraction 0.983). (b) CP with periodic boundary conditions (packing fraction 0.967). In both cases the ratio between the largest and smallest radii is 1575. One can notice some inhomogeneities and voids in the configuration generated with the CP protocol, see the discussion in Sec. IV.1.
We consider a set of $N$ disks obeying a power-law distribution with the exponent $D$. In two dimensions, the exponent satisfies the condition $20<D<$ 2. The number of disks $dN(r)$ whose radii fall within the range ($r,r+dr$) is proportional to $dr/r^{D+1}$. The radii vary within the range of the distribution, from $a$ to $R$. The size ratio $R/a$ determines the length of the range of the distribution on a logarithmic scale. For a finite number of disks, a choice of specific radii that follow the power-law distribution was described in detail in Refs. [6, 8].
RSA and DT protocols enable a compact packing of the disks within a finite area, the exponent of the distribution should be restricted to $D_{\mathrm{Ap}}<D<2$ with $D_{\mathrm{Ap}}=1.3057\ldots$ being the dimension of the Apollonian packing (see the discussion in Ref. [6]). The disks are arranged into a square by placing them from the largest to the smallest. As the set of already placed disks grows, checking their collisions takes more and more computational time. The DT algorithm reduces computation costs of searching for an empty space to add a new disk, which makes it especially effective at high packing fractions. The total number of operations for dense packing with and without DT is $O(N\log N)$ and $O(N^{2})$, respectively (see details of the DT protocol in our paper [8]). We employ the code from the CGAL software package [16].
@@
Figure 2: The smeared structure factor (a) and mass-radius relation (b) for different protocols when the ratio between the largest and smallest radii is 292. The number of particles is $N=125000$ for each protocol. The results for the DT and CP protocols are represented in solid black and blue lines, respectively. The dashed orange lines show the combination of the DT and CP protocols (see Sec. IV.1), the corresponding curves are shifted vertically by a factor of 10 for better visualization. Vertical dotted lines denote the borders of the range over which the fit has been performed. Red curves represent errors after averaging over 20 trials.
III Dense random packings
@@
Figure 3: The smeared structure factor (a) and mass-radius relation (b) for different protocols when the ratio between the largest and smallest radii is 1575. The results for the boundary-trimmed construction (see Fig. 5) is shown in solid green lines, and for the DT and CP protocols in solid black and blue lines, respectively. The corresponding curves are shifted vertically by a factor of 10 for better visualization. Vertical dotted lines denote the borders of the range over which the fit has been performed. Red curves represent errors after averaging over 20 trials.
@@
Figure 4: The smeared structure factor (a) and mass-radius relation (b) for DT protocol with periodic boundary conditions when the ratio between the largest and smallest radii is 2500. Notations are the same as in Fig. 2.
@@
To test the consistency of the fractal properties of dense random packings, they are generated using the DT protocol for different size ratios $R/a$ and boundary conditions. Figures 2, 3, and 4 show the structure factor (2) and mass-radius relation (1) for the total number of disks $N=125000$ and the ratios $292$, $1575$, and $2500$, respectively. Each value of $S(q)$ and $M(r)$ is obtained numerically by averaging over 20 generated trails. The structure factor is smeared with a lognormal distribution function, as described in Ref. [6]. This procedure allows us to eliminate numerous minima and maxima that are superimposed on the decay of the structure factor. The fractal ranges $2\pi/R\ll q\ll 2\pi/a$ and $a\ll r\ll R$ for $S(q)$ and $M(r)$, respectively, increase with increasing the size ratio.
@@
Our analysis reveals (see Table 1) that the fractal properties exhibit a high degree of consistency even for moderate values of the size ratio. The exponents of the mass-radius relation $D_{\mathrm{f}}$, the structure factor $\alpha$, and size distribution $D$ coincide very closely in accordance with the model of dense random packing developed in the previous papers [6] and [8]. The exception is the exponents for $R/a=1575$ and zero boundary conditions (see Fig. 3). They slightly deviate form each other and $D=1.5$. We explain this behavior by the influence of zero boundary conditions: the boundaries of the square “attract” disks of small radii and thus make the configuration slightly inhomogeneous. To reduce the influence of the boundary effects, a part of the disks in contact with the borders of the square is removed (see Fig. 5), and the structure factor and mass-radius relation are calculated for the remaining part of the system, see Fig.  3. All three exponents for the ”boundary-trimmed” set of disks become very close again.
@@
Table 1: The power-law exponents for the structure factors $\alpha$ [$S(q)\sim 1/q^{\alpha}$] and mass-radius relation $D_{\mathrm{f}}$ [$M(r)\sim r^{D_{\mathrm{f}}}$] shown in Figs. 2-4. The exponents are obtained by linear fits within the fractal ranges $Rq_{\text{min}}\leqslant Rq\leqslant Rq_{\text{max}}$ and $r_{\text{min}}/R\leqslant r/R\leqslant r_{\text{max}}/R$, respectively. BC and $\phi$ denotes the boundary conditions and packing fraction, respectively.
Protocol BC N ${\frac{R}{a}}$ $\phi$ $\alpha$ $Rq_{\text{min}}$ $Rq_{\text{max}}$ $D_{\mathrm{f}}$ ${\frac{r_{\text{min}}}{R}}\times 10^{-3}$ ${\frac{r_{\text{max}}}{R}}\times 10^{-3}$ Ref.
DT periodic 125000 292 0.955 1.52 $\pm$ 0.02 5.79 65 1.501 $\pm$ 0.001 9.09 110 Fig. 2
DT(0.95) + CP periodic 125000 292 0.966 1.49 $\pm$ 0.03 5.07 30.41 1.501 $\pm$ 0.003 19.9 145 Fig. 2
CP periodic 125000 292 0.953 1.31 $\pm$ 0.01 6.74 114.65 1.419 $\pm$ 0.002 10.1 148 Fig. 2
DT zero 125000 1575 0.983 1.47 $\pm$ 0.01 5.12 107.12 1.514 $\pm$ 0.001 14.0 177 Fig. 3
DT zero 85000absent85000$\approx$ 85000≈ 85000222With approximately 40000 points near the square borders being removed to avoid the influence of the boundary conditions, see Fig. 5. 1575 0.983 1.50 $\pm$ 0.02 6.94 99.28 1.497 $\pm$ 0.004 14.0 101 Fig. 3
CP periodic 125000 1575 0.967 1.41 $\pm$ 0.01 5.12 310 1.441 $\pm$ 0.004 3.64 177 Fig. 3
DT periodic 125000 2500 0.987 1.50 $\pm$ 0.02 5.96 197 1.502 $\pm$ 0.004 4.22 121 Fig. 4
IV Packing with constant pressure protocol
In this section, we study the characteristics of jammed packings generated through the CP protocol for a square with imposed periodic boundary conditions. The protocol was described in Sec. II above.
IV.1 Moderate size ratio
We realize numerically the CP protocol of dense packing with the parameters $D=1.5$ and $R/a=292$ for 20 trials. The structure factor and mass-radius relation are shown in Fig. 2. The obtained fractal exponents of mass-radius relation $D_{\mathrm{f}}=1.419\pm 0.002$ and structure factor $\alpha=1.31\pm 0.01$ are inconsistent and significantly differ from the power-law exponent $D=1.5$, see Table 1. The observed discrepancies can be explained by the insufficient value of the size ratio, which is discussed in more detail in Sec. IV.2 below. Note that the chosen size ratio $R/a$ even exceeds the $200$ value considered in the paper [7], for which the fractal exponent $\alpha=1.28\pm 0.01$ was numerically obtained at the same $D$. This value is quite close to our result.
To test how jamming itself affects the features of dense packings, a combination of the DT and CP protocol is applied. At the beginning, the system is packed with the DT protocol to the higher packing fraction $0.955$ and after that the CP protocol is realized. The results for the combination of the packings and the single non-jammed DT packing are shown in Fig. 2. For both DT+CP and DT protocols, the fractal and power-law exponents agree very closely, see Table 1. It follows from these findings that jamming itself is not a significant factor influencing the fractal behaviour of the system.
Another simple argument can be made to show the insignificance of jamming per se. Suppose we have a perfectly jammed system of packed disks. Let us scale the system with the factor $1+\varepsilon$ but leave the radii of the disks unchanged. It is clear that for arbitrary small $\varepsilon>0$ no disk is in contact with the others anymore, and then the fraction of rattlers is equal to one. On the other hand, the correlation properties, including the structure factor and mass-radius relation, remain practically unchanged.
As seen in Fig. 1b, the packing obtained using the CP protocol contains empty cavities filled with a few rattlers. Such cavities are formed when several large particles come into contact, forming an almost empty space between them into which smaller particles can no longer penetrate. We believe that it is the cavities and inhomogeneities that are responsible for the non-randomness of the CP packing. In turn, non-randomness is responsible for the inconsistency of the fractal exponents.
IV.2 High size ratio
To investigate the influence of the size ratio on the fractal exponents, the CP packings were obtained with high size ratio $R/a=1575$ for 20 trails, see the Fig. 3 and Table 1 for the results. The fractal exponents of mass-radius $D_{\mathrm{f}}=1.441\pm 0.004$ and structure factor $\alpha=1.41\pm 0.01$ differ from that of obtained for $R/a=292$ and are still inconsistent. However, their values approach each other and to the power-law exponent. We conclude that the CP protocol yields inconsistent fractal behaviour that depends on the initial density and particle size ratio, which leads to ambiguities in its interpretation.
Nevertheless, these results suggest that both fractal exponents $D_{\mathrm{f}}$ and $\alpha$ slowly converge to the power-law exponent $D$ as the size ratio $R/a$ increases. The slow convergence appears to be a result of non-randomness of the CP packings, which becomes less important at high size ratios.
However, the CP computational cost grows dramatically as the size ratio increases, which presents a practical obstacle to achieving full convergence in our modeling framework. A large ratio of the heaviest-to-lightest masses of the particles leads to an apparently high difference in their average velocities, which increases the number of time steps to reach equilibrium. Let us estimate the number of steps in simulations. We assume that particle velocity is approximately proportional to the inverse square root of its mass, which is equal, up to a factor, to the inverse radius in two dimensions. Then we have for the highest and lowest velocities: $v_{\mathrm{max}}\sim 1/a$ and $v_{\mathrm{min}}\sim 1/R$, respectively. The time step $\tau$ should be less than $a/v_{\mathrm{max}}\sim a^{2}$ so that the small particles have time to interact and not fly through each other. The total simulation time should be at least $t_{\mathrm{sim}}\sim R/v_{\mathrm{min}}\sim R^{2}$, so that during the simulation time the largest particle has time to travel a distance not less than its radius. Then the required number of steps scales as the square of the size ratio: $t_{\mathrm{sim}}/\tau\sim R^{2}/a^{2}$. Another limitation is due to the large number of neighboring particles, which reduces the efficiency of LAMMPS parallel computation, because computational resources are mainly spent on data transfer between message-passing-interface threads.
V Conclusions
The model of dense packing developed in previous publications [6, 8] states that all three exponents $D_{\mathrm{f}}$, $\alpha$, and $D$ (for the mass-radius relation, structure factor, and power-law distribution, respectively) coincide if the packing is random, the density is high, and the size ratio $R/a$ is sufficiently large. Deviations from this behavior indicate a violation of at least one of the above conditions.
The results obtained agree with this statement. The DT and RSA protocols are random by construction, and if the other conditions are fulfilled then indeed we have $D_{\mathrm{f}}=\alpha=D$ within computational error (see Sec. III). The CP protocol is apparently non-random as one can see from Fig. 1, while jamming per se has no significant effect on the fractal properties (see Sec. IV.1). We believe that non-randomness of CP packings is the main reason of inconsistency of the resulting fractal properties. Since $D_{\mathrm{f}}\not=\alpha$ and the exponents depend on the initial packing conditions and the size ratio (see Sec. IV), they lack universality, and it is not clear how to describe them theoretically at all.
Nevertheless, our findings suggest that if a packing in two or three dimensions is even non-random then the exponents $D_{\mathrm{f}}$ and $\alpha$ slowly converge to $D$ in the limit $R/a\to\infty$ and the packing fraction tending to one. This is a stronger statement than was suggested previously [6, 8]. The question then arises how to describe randomness quantitatively for a power-law distribution of radii and how to evaluate its effect on the inconsistency of fractal properties for a finite packing. This is an unsolved problem that opens new research perspectives.
Appendix A General definitions and relations
In this appendix we follow Section II of our previous paper [6], where the definitions and relations are explained in more detail.
For a set of $N$ points of unit weight located at the positions $\bm{r}_{1},\cdots,\bm{r}_{N}$, the mass-radius relation $M(r)$ is defined as the average value of mass enclosed in the imaginary circle of radius $r$, which is centered on a point belonging to the set [13]. According to the definition, it is given by
$M(r)=\frac{1}{N}\sum_{i,j}\theta(r-r_{ij})=1+\frac{1}{N}\sum_{i\neq j}\theta(r-r_{ij}),$ (1)
where $r_{ij}=|\bm{r}_{i}-\bm{r}_{j}|$ and $\theta(z)$ is the Heaviside step function, that is, $\theta(x)=1$ for $x\geqslant 0$ and zero elsewhere. Then $M(r)=1$ when $r$ is less than the smallest distance between points and $M(r)=N$ when $r$ exceeds the largest distance.
The structure factor of the set of points is defined as [19]
$S(q)=\frac{1}{N}\left\langle\rho_{\bm{q}}\rho_{-\bm{q}}\right\rangle_{\hat{q}},$ (2)
where $\rho_{\bm{q}}=\sum_{j}e^{-i\bm{q}\cdot\bm{r}_{j}}$ is the Fourier transform of the density of the points $\rho(\bm{r})=\sum_{j}\delta(\bm{r}-\bm{r}_{j})$, and the brackets $\langle\cdots\rangle_{\hat{q}}$ stand for the average over all directions of unit vector $\hat{q}$ along $\bm{q}$. By definition, the structure factor depends only on the absolute value of $\bm{q}$, which we denote as $q$. It obeys the conditions $S(q)\simeq 1$ when $q\to\infty$ and $S(q)=N$ at $q=0$. The structure factor can be measured in small-angle scattering experiments [20].
In two dimensions, mass radius and structure factor are related by the following equation
$\frac{1}{r}\frac{\partial M}{\partial r}=\,$ $\frac{1}{2\pi}\int d^{2}\!q\,e^{i\bm{q}\cdot\bm{r}}[S(q)-1]$
[…]
where $J_{0}(z)$ is the Bessel function of zeroth order. When $S(q)-1\sim 1/q^{D_{\mathrm{f}}}$ for $q\to\infty$ and the exponent lies within the range $21<D_{\mathrm{f}}<$ 2, then its two-dimensional Fourier transform $\frac{1}{r}\frac{\partial M}{\partial r}$ is proportional [14] to $1/r^{2-D_{\mathrm{f}}}$ for $r\to 0$. It follows that $M(r)\sim r^{D_{\mathrm{f}}}$ at sufficiently small values of $r$. Note that for the densely packed set of disks described in Sec. II, such infinite-range asymptotics of the structure factor can be realized only in the limit of the infinite size ratio: $R/a\to\infty$.

### v2

# The influence of packing protocol, size ratio, and pore structure on fractal exponents in dense polydisperse packings

@@ Abstract
We study fractal properties of systems of densely and randomly packed disks, obeying a power-law distribution of radii, which is generated by using various protocols: Delaunay triangulation (DT) and constant pressure (CP) protocols, and the generalized Apollonian packing. The power-law exponents of the mass-radius relation and structure factor are obtained numerically for various values of the size ratio of the distribution, defined as the largest-to-smallest radius ratio. We show that the size ratio is an important control parameter responsible for the consistency of the fractal properties of the system: the greater the ratio, the less the finite size effects are pronounced and the better the agreement between the exponents. For the DT protocol, all three exponents coincide even at moderate values of the size ratio. For the CP protocol, the exponents are different for both moderate and large size ratios. The suppression of the exponent of the structure factor in the CP packing is explained by the specific behaviour of pores, which contain relatively big cavities. We develop an algorithm for calculating the pore size distribution and show that it is related to the exponent of the structure factor. We argue that the presence of the cavities lowers the configuration entropy and thus reduces the randomness of the CP packing. Thus the cavities reduce both packing fraction and randomness of the CP packings. Nevertheless, there is a tendency for the exponents to converge as the size ratio increases, suggesting that all the exponents become equal in the limit of infinite size ratio.
dense random packings; jammed packings; constant pressure protocol; generalized Apollonian packing; pore size analysis; power-law polydispersity; fractals; structure factor; mass-radius relation.
@@ I Introduction
Compact packings with power-law size distributions are of great interest for various applications in science and engineering, in particular for understanding the structural properties of materials [1]. This understanding is critical for applications ranging from ultra-high performance concrete [2] to biological systems [3] and soils [4], as it enables the design and optimization of materials with desired mechanical properties and stability.
Analytical solutions were obtained for the jamming fractions of polydisperse packings of spheres [5] and disks [6] in the case of moderate variations of their sizes (see also Ref. [7]). The case of a high size ratio of the power-law polydispersity was considered in the papers [8, 9, 10, 11]. The spatial correlations in such systems were found to exhibit fractal-like behaviour. The description of the emerging fractal properties is one of the most interesting and intriguing problems. In this paper, we continue to explore these issues.
In numerical simulations, dense packing can be achieved by various methods (see [1] for a review), including random successive addition (RSA) [12, 13]. In RSA, particles are added one at a time in random positions. If a particle overlaps with existing ones, it is discarded. This process continues until no more particles can be added without overlap. Recently, the authors suggested [11] Delaunay triangulation protocol as a modification of RSA. DT creates a network of triangles connecting the centers of packed particles, which optimizes the process of finding empty space for newly added particles.
Another method of dense packings is the constant pressure protocol, which uses LAMMPS (large-scale atomic/molecular massively parallel simulator) with GRANULAR package [14]. It allows the modeling of granular materials under controlled pressure conditions. Constant external pressure slowly compresses an initially dilute configuration of particles, leading to a compact jammed packing.
We also consider the generalized Apollonian packing (GAP) [15], which is a deterministic construction that extends the classical Apollonian packing [16] and enables us to get the fractal exponents up to $1.803\ldots$. Unlike the DT and CP protocols, GAP does not pack a given set of disks, but generates it deterministically. The GAP is included in this paper mainly as a reference for pore size analysis (see Sec. VI below) but not as a random packing protocol. We consider a construction of GAP, which yield the exponent $1.46$ being quite close to $D=1.5$ (see Sec. III below).
It was shown with RSA [9] and DT [11] protocols that dense random packings of disks with a power-law size distribution exhibit fractal-like properties with the fractal dimension $D_{\mathrm{f}}$ being equal to the exponent of the power-law size distribution $D$. On the other hand, Monti et al. [10] obtained jammed packings of the same distribution with CP protocol and argued that the packings, although exhibiting fractal properties, nevertheless have fractal dimensions different from the exponent of the power-law size distribution. The authors of the paper [10] concluded that the fractal exponent depends on packing protocol.
@@
As shown in the previous paper [9], the fractal range of densely packed disks or spheres with a power-law size distribution is determined by their largest-to-smallest size ratio, and this result is still valid in the thermodynamic limit [11]. In the paper [10], the ratio was taken in two dimensions from $100$ to $300$ for various exponents $D$ of the distribution [^1] Note that in the paper [10], the notation $\beta=D+1$ is used.. It is shown below in Sec. V that these ratios are not large enough to obtain the correct value of the fractal dimension of the system. To this end, the data of jammed packings with CP protocol of Ref. [10] are reproduced for a specific value of the exponent $D=1.5$ and the size ratio $292$. We calculate both structure factor and mass-radius relation and obtain the discrepancy between the exponents $\alpha$ $S(q)\sim 1/q^{\alpha}$] and $D_{\mathrm{f}}$ $M(r)\sim r^{D_{\mathrm{f}}}$] in the corresponding fractal ranges: $\alpha\simeq 1.31$ and $D_{\mathrm{f}}\simeq 1.419$, see Fig. 3 and Table 1 below. This discrepancy is due to the insufficient size ratio.
Increasing the size ratio to $R/a=1575$ and keeping the other parameters unchanged (see Sec. V.2 below), we obtain $\alpha\simeq 1.41$ and ${D_{\mathrm{f}}}\simeq 1.441$, which are much closer to each other, although its value is still different from $D=1.5$ predicted by our model of dense random packings.
To understand the finite-size effects in CP packing, the pore size distribution is studied. We have developed an algorithm that effectively represent pores as a set of disks (see Sec. VI below). It is shown that the CP protocol generates relatively large cavities, which affect the fractal exponents and decrease their values. The pore size distribution is strongly influenced by the size ratio of the particle distribution. Random packings generated by the DT protocol satisfy $D_{f}=\alpha=D$ once the size ratio is sufficiently large, whereas CP packings exhibit some deviations due to the formation of large cavities that reduce randomness and modify the effective pore-size distribution. Nevertheless, as the size ratio increases, the influence of the cavities diminishes, and the fractal exponents obtained with different protocols show a clear tendency to converge to $D$ (see Table 1 below).
Thus, although different packing protocols lead to clearly different configurations, the apparent dependence of fractal exponents on the protocol is not fundamental. Instead, the decisive factor determining the fractal properties of dense polydisperse packings is the degree of configurational randomness, which is closely related to the pore statistics.
This paper is organized as follows. In Sec. II, we explain in detail a set of disk to be packed and packing methods. In the next section, the GAP is considered and the parameters of its generation are described. The fractal properties of dense random packing with DT protocol are studied for various values of the size ratio in Sec. IV. In Sec. V, the results of packing with CP protocol are represented. In the next section, we find the pore size distribution in various protocols and examine its effect on the fractal exponent of the structural factor. In Conclusions, we summarize the main results and outline prospects for future research.
@@ II A set of disks to be packed and methods
Refer to captionRefer to captionRefer to caption
Figure 1: Comparison of spatial distributions of $N=125000$ disks obtained from different protocols: (a) DT with zero boundary conditions (packing fraction 0.983). (b) CP with periodic boundary conditions (packing fraction 0.967). The disk borders and pores are highlighted in orange and white, respectively. One can notice some inhomogeneities and voids in the configuration generated with the CP protocol, see the discussion in Sec. V.1 below. (c) GAP (packing fraction 0.943) and N = 52366. In all cases the ratio between the largest and smallest radii is 1575.
Let us consider a set of $N$ disks obeying a power-law distribution with the exponent $D$. We introduce a cumulative distribution and denote the number of disks, whose radii are bigger than or equal to $r$, by $N(r)$. Then we have $N(r)\sim 1/r^{D}$. In two dimensions, the exponent satisfies the condition $0<D<2$. The number of disks $|dN(r)|$ whose radii fall within the range ($r,r+dr$) is proportional to $dr/r^{D+1}$. The radii vary within the range of the distribution, from $a$ to $R$. The size ratio $R/a$ determines the length of the range of the distribution on a logarithmic scale. For a finite number of disks, a choice of specific radii that follow the power-law distribution was described in detail in Refs. [9, 11].
RSA and DT protocols enable a compact packing of the disks within a finite area, the exponent of the distribution should be restricted to $D_{\mathrm{Ap}}<D<2$ with $D_{\mathrm{Ap}}=1.3057\ldots$ being the dimension of the Apollonian packing [20] (see the discussion in Refs. [9]). The disks are arranged into a square by placing them from the largest to the smallest. As the set of already placed disks grows, checking their collisions takes more and more computational time. The DT algorithm reduces computation costs of searching for an empty space to add a new disk, which makes it especially effective at high packing fractions. The total number of operations for dense packing with and without DT is $O(N\log N)$ and $O(N^{2})$, respectively (see details of the DT protocol in our paper [11]). We employ the code from the CGAL software package [21].
@@
Figure 2: The smoothed structure factor (a) and mass-radius relation (b) for GAP when the ratio between the largest and smallest radii is 1575. Vertical dotted lines denote the borders of the range over which the fit has been performed
III Generalized Apollonian Packing
In addition to the random packing protocols discussed above, we also employ a deterministic hierarchical construction based on the generalized Apollonian packing also known as a space-filling bearing. This deterministic structure provides a useful benchmark configuration, against which the statistical properties of the random packings generated by the DT and CP protocols can be compared.
We follow the algorithm of Oron and Herrmann [15], which enables us to construct the sets with the fractal dimension ranging approximately from $1.306$ to $1.803$. The method generates a strip geometry that is invariant under a prescribed set of Möbius transformations. The transformations are reflections and inversions chosen such that the resulting configuration contains no frustrated loops and completely fills the strip in the limit of infinite number of iterations. Below we describe the parameters of the algorithm that we used and refer the reader to the paper [15] for details.
The construction begins from a “base loop” consisting of $l$ circles, which are mapped onto themselves under two inversions $I_{1}$ and $I_{2}$ centered at the outermost contact points of the seed. For the F2 family used in this work, the seed is symmetric with respect to two vertical reflection axes, and the alignment of the contact points ensures that the inversion-circle centers lie on a line orthogonal to the strip boundaries. The parameters $n_{1}$ and $n_{2}$ determine the radii of the inversion circles through the relations $r_{1}^{2}=a^{2}z_{n_{1}}$ and $r_{2}^{2}=a^{2}z_{n_{2}}$, with $z_{k}=\left[\cos\!\left(\pi/(k+3)\right)\right]^{-2}$, and $a=(z_{n_{1}}+z_{n_{2}})^{-1/2}$ [see Eq. (A.8) in Ref. [15]] denoting half of the strip period. Repeated application of the transformation sequences $R_{1}I_{1},(R_{1}I_{1})^{2},\cdots,(R_{1}I_{1})^{n_{1}+2}$ and $R_{2}I_{2},(R_{2}I_{2})^{2},\cdots,(R_{2}I_{2})^{n_{2}+2}$ generates an iterative cascade of circles whose radii decrease geometrically, thereby completely filling the available space. Here, $R_{1}$ and $R_{2}$ are reflections with respect to vertical lines at 0 and $a$, respectively.
In this study, we use the simplest member of the F2 family with parameters $l=4$, $n_{1}=0$, and $n_{2}=0$, corresponding to the top-left structure in Fig. 7 of Ref. [15]. According to the analysis reported in Fig. 10 of the same reference, this configuration has the fractal dimension $D_{\mathrm{f}}\approx 1.46$, placing it near the lower end of the spectrum of space-filling bearing geometries. The structure factor and mass-radius relation are shown in Fig. 2. For smoothing we used [9] a log-normal distribution with the relative variance $\sigma_{\mathrm{r}}=0.25$. This procedure allows us to eliminate numerous minima and maxima that are superimposed on the decay of the structure factor.
Note that the usual Apollonian packing corresponds [15] to the limiting case of this generalized construction with the smallest base–loop size $l=4$ and infinite inversion depth, $(n_{1},n_{2})\to(\infty,\infty)$. In this regime the inversion circles have radii $r_{1}=r_{2}=a$, and the iterative Möbius transformations reproduce the classical Apollonian gasket [16] with the fractal dimension $D_{\mathrm{f}}=1.306\ldots$.
Figure 3: The smoothed structure factor (a) and mass-radius relation (b) for different protocols when the ratio between the largest and smallest radii is 292. The number of particles is $N=125000$ for each protocol. The results for the DT and CP protocols are represented in solid black and blue lines, respectively. The dashed orange lines show the combination of the DT and CP protocols (see Sec. V.1), the corresponding curves are shifted vertically by a factor of 10 for better visualization. Vertical dotted lines denote the borders of the range over which the fit has been performed. Red curves represent errors after averaging over 20 trials.
IV Dense random packings with DT protocol
Figure 4: The smoothed structure factor (a) and mass-radius relation (b) for different protocols when the ratio between the largest and smallest radii is 1575. The results for the boundary-trimmed construction (see Fig. 6) is shown in solid green lines, and for the DT and CP protocols in solid black and blue lines, respectively. The corresponding curves are shifted vertically by a factor of 10 for better visualization. Vertical dotted lines denote the borders of the range over which the fit has been performed. Red curves represent errors after averaging over 20 trials.
@@ IV Dense random packings with DT protocol
Figure 5: The smoothed structure factor (a) and mass-radius relation (b) for DT protocol with periodic boundary conditions when the ratio between the largest and smallest radii is 2500. Notations are the same as in Fig. 3.
@@
To test the consistency of the fractal properties of dense random packings, they are generated using the DT protocol for different size ratios $R/a$ and boundary conditions. Figures 3, 4, and 5 show the structure factor (3) and mass-radius relation (2) for the total number of disks $N=125000$ and the ratios $292$, $1575$, and $2500$, respectively. Each value of $S(q)$ and $M(r)$ is obtained numerically by averaging over 20 generated trails. The smoothing function is a log-normal distribution with the relative variance $\sigma_{\mathrm{r}}=0.2$. The fractal ranges $2\pi/R\ll q\ll 2\pi/a$ and $a\ll r\ll R$ for $S(q)$ and $M(r)$, respectively, increase with increasing the size ratio.
@@
Our analysis reveals (see Table 1) that the fractal properties exhibit a high degree of consistency even for moderate values of the size ratio. The exponents of the mass-radius relation $D_{\mathrm{f}}$, the structure factor $\alpha$, and size distribution $D$ coincide very closely in accordance with the model of dense random packing developed in the previous papers [9] and [11]. The exception is the exponents for $R/a=1575$ and zero boundary conditions (see Fig. 4). They slightly deviate form each other and $D=1.5$. We explain this behaviour by the influence of zero boundary conditions: the boundaries of the square “attract” disks of small radii and thus make the configuration slightly inhomogeneous. To reduce the influence of the boundary effects, a part of the disks in contact with the borders of the square is removed (see Fig. 6), and the structure factor and mass-radius relation are calculated for the remaining part of the system, see Fig.  4. All three exponents for the ”boundary-trimmed” set of disks become very close again.
@@
Table 1: The power-law exponents for the structure factors $\alpha$ $S(q)\sim 1/q^{\alpha}$] and mass-radius relation $D_{\mathrm{f}}$ $M(r)\sim r^{D_{\mathrm{f}}}$] shown in Figs. 3-5. The exponents are obtained by linear fits within the fractal ranges $Rq_{\text{min}}\leqslant Rq\leqslant Rq_{\text{max}}$ and $r_{\text{min}}/R\leqslant r/R\leqslant r_{\text{max}}/R$, respectively. BC and $\phi$ denotes the boundary conditions and packing fraction, respectively.
Protocol BC N ${\frac{R}{a}}$ $\phi$ $\alpha$ $Rq_{\text{min}}$ $Rq_{\text{max}}$ $D_{\mathrm{f}}$ ${\frac{r_{\text{min}}}{R}}\times 10^{-3}$ ${\frac{r_{\text{max}}}{R}}\times 10^{-3}$